# Balancing Specialization and Adaptation in a Transforming Scientific Landscape

### 03 Jan 2024 v2

@@ Abstract
How scientists navigate between the need to capitalize on their prior knowledge by specializing, and the urge to adapt to evolving research opportunities? Drawing from diverse perspectives on adaptation, in particular from institutional change and cultural evolution, this paper proposes a Bayesian model of the evolution of scientists’ research portfolios in response to transformations in their field. The model relies on scientific abstracts and authorship data to evaluate the influence of intellectual, social, and institutional resources on scientists’ trajectories within a cohort of 2 19521952\,1952 195 high-energy physicists between 2000 and 2019. The reallocation of research efforts is shown to be structured by learning costs, thus enhancing the utility of the scientific capital disseminated among scientists. Two dimensions of social capital, namely “diversity” and “power”, have opposite effects on the magnitude of change in scientists’ research interests: while “diversity” disrupts and expands research interests, “power” stabilizes physicists’ research agendas – as does institutional stability. Social capital plays a more crucial role in shifts between cognitively distant research areas.
@@
adaptation; specialization; cultural evolution; science of science; human mobility; optimal transport
@@ 1 Introduction
Scientists are subject to conflicting incentives. On the one hand, they must work within the realm of their expertise, where they can most effectively exploit their prior knowledge and compete with peers; this conservative preference for familiar research topics is at the root of specialization. On the other hand, scientists are simultaneously compelled to revise their research agenda to engage with more promising research areas in order to benefit from more exposure, to secure funding, or to seize job opportunities. Thus, in some instances, specialization is at odds with the need to adapt to the decline of certain research opportunities and the growth of new ones. How do scientists navigate the trade-off between specialization (i.e. the concentration of their intellectual resources within a narrow cognitive range) and adaptation (i.e. the need to adjust these resources to new realities)? This conflict differs from the “essential tension” between “tradition” and “innovation” proposed by [1, ], or that between “exploration” and “exploitation” [2], which have both been explored quantitatively in previous works [3, 4, 5, 6, 7, 8, 9]. First, “adaptation” is not tantamount to innovation or disruption, for it can be a conformist move (e.g. as a result of a bandwagon effect, [10]). Moreover, unlike “exploration”, adaptation is not a search strategy or a method for acquiring information, but rather the convergence towards a new state more congruent with current realities. Disruptions due to breakthroughs in Machine Learning (e.g. generative Large Language Models) or challenges due to climate change make it particularly pressing to understand the factors that enhance scientists’ adaptability to changing circumstances. Therefore, the present paper investigates scientists’ responses to changes in their field (whether it is driven by epistemic or institutional factors), and the effect of their capital (intellectual, social, or institutional) on their ability to adapt. Drawing insights on adaptation from cultural evolution and institutional change, we develop an unsupervised Bayesian approach to analyze changes in scientists’ research agenda while measuring the effect of “capital” (intellectual or social) on their individual trajectories. The model is applied to a cohort of high-energy physicists between the years 2000 and 2020, a time during which the historical driver of progress in the field – particle accelerators – have been contested by emerging astrophysical experiments, thus transforming the landscape of opportunities.
@@
Our approach reveals trends in the field: the boom of dark matter research – fueled by migrations away from the physics of neutrinos and the electroweak sector – and the partial disintegration of string theory into the study of black holes and holographic dualities. More importantly, this analysis also shows that changes in scientists’ research portfolios are structured by learning costs, as scientific communities adapting to new circumstances address an Optimal Transport problem by reallocating research efforts efficiently. Moreover, two dimensions of social capital, namely “diversity” and “power” [11], have opposite effects on change. While “diversity” of social capital (i.e. the extent to which scientists have access to diverse cognitive resources via their collaborators) promotes change and further diversification of scientists’ research interests, “power” (the size of their network) stabilizes their research interests. Social capital has a stronger effect on transfers between research areas that are more cognitively “distant”. Furthermore, institutional stability is associated with more conservative research interests. Overall, we contribute: i) a theoretical account of the features of change in scientists’ research interests; ii) a novel methodological approach that introduces a model of scientists’ trajectories and measures of intellectual capital, social capital, diversity, and power; and finally, iii) empirical evidence from High-Energy Physics (HEP). More generally, this paper addresses the relative lack of empirical works within the body of literature that investigates science as a cultural evolutionary system [12].
@@ 1.2 Conceptual framework
Scientists’ opportunities are constrained by their prior knowledge, since entering new research areas involves learning costs. Yet, the exploitation of prior knowledge may require adjustments to new realities. Through a gradual process of adaptation, scientists can strategically retain the benefits of “problem retention” (e.g., the exploitation of “accumulated skills and resources” in one area, or “of an established research network”, [13, p. 106]) while progressively investing resources in new research directions. This is illustrated in Figure 1, which represents the research agendas of one scientist during two distinct time periods. Cells indicate the resources exploited by the scientist (e.g., concepts, models, methods, or techniques) and colors indicate to what problem areas this knowledge applied. A resource can be applied simultaneously to different ends, as expressed by the hatched cell; however, resources are often tailored for specific research purposes. The figure also shows how scientists can enter new research areas by repurposing their knowledge to new ends [18, 19]. We call this strategy “conversion”, in reference to the typology of incremental institutional change proposed by [20, ]. Indeed, as shown in previous works on the transformations of high-energy physics facilities to photon science instruments [21, 22, 23], historical institutionalism can account for gradual adaptations with large cumulative effects taking place in response to scientific and technological change [24]. In this paper, we apply the typology of change to individuals rather than organizations.
@@
Scientists manage two kinds of assets when navigating the trade-off between specialization and adaptation: their own prior expertise, and the expertise to which they have access through their social network. We propose to consider both of these as “capital” [17], which designates the different assets that individuals (e.g. scientists) can accumulate and leverage in the competitive context of their field. “Capital” (whether “economic”, “cultural”, “social” or even “symbolic”, cf. [17]) defines the scope of scientists’ opportunities and therefore their ability to revise their research agenda. This paper considers the cultural capital possessed by scientists in the form of scientific knowledge (which will be referred to as “intellectual capital”), and social capital. Measures that represent these concepts will be proposed, and their effect on the magnitude of transfers of attention across research areas will be evaluated. Emphasis will be put on the divide underlined in [11] between two dimensions of social capital, namely “power” (in fact, network size in the present paper) and “diversity” (of cognitive resources). Group diversity is generally recognized as a factor of adaptation in an evolving environment [25, 26, 27] or in the context of problem solving [28]. “Power” is also plausibly associated with higher abilities.
@@ 1.3 The case of high-energy physics: navigating a changing epistemic landscape
The field of particle physics has relied on the input of increasingly large particle colliders to achieve progress. These efforts culminated in 2010 with the start of the Large Hadron Collider (LHC), the largest accelerator ever. However, the LHC has found no evidence for anything that was not already predicted by the Standard Model of particle physics, and it is increasingly plausible that no future accelerator could ever find any evidence for new particles, leading to a situation of “crisis” [29]. Although the LHC will continue to take data for years and plans for future accelerators are being discussed [30], some have even speculated that particle physics as we know it has come to an end [31, 32], “the proscenium [being] captured by astrophysics and cosmology” instead [32]. In a recent assessment of the field, physicist Mikhail Shifman claimed that the “pause in accelerator programs we are witnessing now is not necessarily […] the end of explorations at [high energies]”; instead, such explorations “will continue, perhaps in a new form, with novel devices” [33]. Indeed, new experimental opportunities have emerged in parallel. Gravitational Waves astronomy began in 2015, with the first “direct” detection of such waves from a black-hole merger by LIGO [34]. Other opportunities include direct searches of dark matter of astrophysical origin in underground facilities (e.g. LUX and Xenon experiments), and more precise data on the early universe from the Planck observatory. In fact, astrophysics seems to be increasingly replacing particle colliders in citations across experimental and theoretical high-energy physics. [35]. These transformations have unfolded gradually: the LHC started in 2010 but it operated much below its target energy until 2015; Planck released data on the cosmic microwave background between 2010 and 2018; etc. (see Figure 9, Appendix A.1). The present paper investigates how high-energy physicists have adapted in reaction to these changes in the scientific landscape.
@@ 2.1 Data
Our source is the Inspire HEP database [36]. It aggregates high-energy physics literature from various sources, including the main scientific publishers and arXiv, and has been used in a number of works [37, 38, 39, 40, 35]. For the literature on HEP, it is more comprehensive than the APS dataset that has been the focus of previous work on the evolution of research interests, because the latter is limited to a few journals [^1] https://journals.aps.org/datasets. The database is free-access and includes semantic data for each paper (mainly their title and abstract, sometimes also keywords and PACS categories), curated authorship and institutional affiliations, lists of references and associated experiments. This database implements both automatized and manual measures for the disambiguation of author names [^2] Besides the use of “advance algorithms” of author-disambiguation, Inspire invites scientists to correct their own publication record on the website (cf. https://twiki.cern.ch/twiki/pub/Inspire/WebHome/INSPIRE_background.pdf, June 2014) , which has previously allowed [39, ] to analyze gender gaps in physicists’ research careers based on these data. Nevertheless, occasional misidentifications remain possible.
@@ 2.2 Measuring research agendas
Our strategy for measuring research portfolios relies neither on PACS categories (in contrast to [4, 14, 5, 7]) nor on citation data (unlike [6]). Instead, research areas are extracted with a topic model that recovers latent “topics” within the corpus and their vocabulary distributions [42, 43]. This makes the approach applicable to a wider range of situations for which only textual data (even short texts) are available. Arguably, texts also provide a more direct access to the structure of knowledge. A coarse-grained classification of the literature is performed into $K=20$ broad “topics”. The number of topics is always somewhat arbitrary, but this is not as much an issue of topic models themselves: as [13, ] puts it, “in such analyses [of problem change], empirical findings will in part reflect the defined scope of problem areas”, which is itself arbitrary. In fact, that topic models give some control over the cognitive scale is a feature, rather than a bug [^4] Previous works based on the PACS categories have leveraged the different levels of this hierarchical classification system to investigate different scales..
We use the Embedding Topic Model developed by [44], which relies on pretrained embeddings representations of the keywords thus providing more reliable classifications of heavy-tailed vocabulary distributions. We train the model the abstracts of the $D=186,162$ documents published between 2000 and 2019. Keywords are extracted from the papers’ titles and abstracts by filtering n-grams between one and three words matching certain syntactic expressions that we can expect to carry scientific information (by designating concepts, models, methods etc.), following the procedure described in [35] and [45]. Keywords’ embeddings are learned using a skip-grap model in a $L=50$ dimension space (few dimensions are needed given the small size of the vocabulary, $V=4,751$). We obtain the topics listed in Appendix A.3, Table 1. This table shows their most frequent keywords, as well as the PACS categories that are most correlated with them. Four of the 20 resulting topics regroup keywords that do not clearly refer to any specific research area (e.g. “paper”, “approach”) and correlate poorly with the PACS categories, which suggests their lack of scientific dimension. Consequently, the present analysis only considers the $K=16$ remaining topics that designate actual research areas (as confirmed by their strong tendency to cite themselves, cf. Appendix A.3.3, and their correlation with the PACS classification, cf. Appendix A.3.4).
@@ 2.3 Measuring capital
Evaluting the effect of intellectual and social capital on scientists’ trajectories requires quantitative measures of these concepts. An extensive review of prior usage of Bourdieu’s notion of capital in bibliometrics has been provided by [46, p. 198-200 ], with most references addressing symbolic and social capital, and none addressing cultural capital. In the following, a consistent approach to intellectual and social capital is proposed, from which measures of diversity and power are derived.
@@ 2.3.2 Social capital
Many measures of scientists’ social capital have been proposed [11, 46], the simplest being the amount of collaborators of a scientist (i.e. degree centrality in the co-authorship network). Other measures are related to betweeness centrality, which concerns the extent to which an actor “bridges” a network (e.g. “brokerage”, i.e., the ability of an individual to overcome “structural holes” in a social network, [47]). The difference between these measures matter; for instance, redundant contacts (and sheer network size) can detrimental in some circumstances. [11, ] distinguishes two general approaches to social capital, depending on whether the emphasis is placed on “power” versus “diversity”. They propose alternative measures based on citations that correlate better with scientists’ performance than measures of degree centrality. All the measures discussed in [11] represent social capital by a single scalar; however, a single number misses multiple dimensions of social capital. In fact, according to [16, ], “the volume of social capital possessed by a particular agent […] depends on the extent of the network of links that he can effectively mobilize, and on the volume of capital (economic, cultural or symbolic) possessed by each of those to whom he is linked”. In that respect, social capital can come in different forms depending on the resources being leveraged via one’s network. In the following, we focus on the intellectual dimension of social capital, which we represent by a vector $\bf{S}_{a}$ that encodes how the amount of expertise in each research area is carried is available to scientists via their network. More specifically, we define $\bf{S}_{a}$ as the sum of the intellectual capital of $a$’s collaborators, weighted by the strength of their relationship:
@@ 2.3.3 Diversity and power
Measures of “diversity” (and “power”) can be readily derived from $\bm{I_{a}}$ (and $\bm{S_{a}}$). A common measure of diversity is the Shannon entropy $H$ [^8] The Shannon entropy of the discrete uniform distribution on $\{1,\dots,n\}$ is $\log{n}$, such as that the exponential of the entropy is roughly the effective amount of outcomes for the distribution ($n$).. Let $D(\bm{I_{a}})=\exp{H(\bm{I_{a}})}$ be the diversity of intellectual capital (and $D(\bm{S_{a}})=\exp{H(\bm{S_{a}})}$ that of social capital). Roughly, these are measures of how many research areas scientists have divided their cognitive or social resources among. In the cohort, individuals typically have cognitive resources in several research areas ($\mu_{D(\bm{I_{a}})}=6.1,\sigma=2.2$), and social capital is even more diverse ($\mu_{D(\bm{S_{a}})}=8.5,\sigma=2.2$). In fact, intuitively, scientific collaborations enable individuals to take advantage of their group’s cognitive diversity. Furthermore, $D(\bm{I_{a}})$ and $D(\bm{S_{a}})$ are highly correlated ($R=0.75$); indeed, individuals with more diverse interests are more able to engage with diverse collaborators. Since the diversity of social capital is only expected to enhance individuals’ abilities when it exceeds that of their own intellectual resources, from now on, only excess social capital diversity $D^{\ast}(\bm{S_{a}})$ (defined as the residuals of the linear regression of $D(\bm{S_{a}})$ against $D(\bm{I_{a}})$, by ordinary least squares) is considered999[7, ] noted the difficulty of determining the direction of the causal relationship between social resources and research interests. This approach aims to overcome the very same challenge..
@@
For each of these (diversity and power), an alternative measure is also considered for robustness assessment (based on Stirling’s index for diversity and brokerage for power). All these measures are compared in Appendix A.4.

### 14 Jan 2024 v3

@@ Abstract
How scientists navigate between the need to capitalize on their prior knowledge by specializing, and the urge to adapt to evolving research opportunities? Drawing from diverse perspectives on adaptation, in particular from institutional change and cultural evolution, this paper proposes an unsupervised Bayesian model of the evolution of scientists’ research portfolios in response to transformations in their field. The model relies on scientific abstracts and authorship data to evaluate the influence of intellectual, social, and institutional resources on scientists’ trajectories within a cohort of 2 19521952\,1952 195 high-energy physicists between 2000 and 2019. The reallocation of research efforts is shown to be shaped by learning costs, thus enhancing the utility of the scientific capital disseminated among scientists. Two dimensions of social capital, namely “diversity” and “power”, have opposite effects on the magnitude of change in scientists’ research interests: while “diversity” disrupts and expands research interests, “power” stabilizes physicists’ research agendas – as does institutional stability. Social capital plays a more crucial role in shifts between cognitively distant research areas. This contribution paves the way for further investigation of science and scientific communities as adaptative systems.
@@
adaptation; specialization; cultural evolution; science of science; computational social science
@@ 1 Introduction
Scientists are subject to conflicting incentives. On the one hand, they must work within the realm of their expertise, where they can most effectively exploit their prior knowledge and compete with peers; this conservative preference for familiar research topics is at the root of specialization. On the other hand, scientists are simultaneously compelled to revise their research agenda to engage with more promising research areas in order to benefit from more exposure or to secure jobs and funding. Thus, in some instances, specialization is at odds with the need to adapt to the decline of certain research opportunities and the growth of new ones. How do scientists navigate the trade-off between specialization (i.e. the concentration of their intellectual resources within a narrow cognitive range) and adaptation (i.e. the need to adjust these resources to new realities)? This conflict differs from the “essential tension” between “tradition” and “innovation” proposed by [1, ], or that between “exploration” and “exploitation” [2], which have both been explored quantitatively in previous works [3, 4, 5, 6, 7, 8, 9]. First, “adaptation” is not tantamount to innovation or disruption, for it can be a conformist move (e.g. as a result of a bandwagon effect, [10]). Moreover, unlike “exploration”, adaptation is not a search strategy or a method for acquiring information, but rather the convergence towards a new state more congruent with current realities. Disruptions due to breakthroughs in Machine Learning (e.g. generative Large Language Models) or challenges due to climate change make it particularly pressing to understand the factors that enhance scientists’ adaptability to changing circumstances. Therefore, the present paper investigates scientists’ responses to changes in their field (whether it is driven by epistemic or institutional factors), and the effect of their capital (intellectual, social, or institutional) on their ability to adapt. Drawing insights on adaptation from cultural evolution and institutional change, we develop an unsupervised Bayesian approach to analyze changes in scientists’ research agenda while measuring the effect of “capital” (intellectual or social) on their individual trajectories. The model is applied to a cohort of high-energy physicists between the years 2000 and 2020, a time during which the historical driver of progress in the field – particle accelerators – have been contested by emerging astrophysical experiments, thus transforming the landscape of opportunities.
@@
Our approach reveals trends in the field: the boom of dark matter research – fueled by migrations away from the physics of neutrinos and the electroweak sector – and the partial disintegration of string theory into the study of black holes and holographic dualities. More importantly, this analysis also shows that changes in scientists’ research portfolios are structured by learning costs, as scientific communities adapting to new circumstances address an Optimal Transport problem by reallocating research efforts efficiently. Moreover, two dimensions of social capital, namely “diversity” and “power” [11], have opposite effects on change. While “diversity” of social capital (i.e. the extent to which scientists have access to diverse cognitive resources via their collaborators) promotes change and further diversification of scientists’ research interests, “power” (the size of their network) stabilizes their research interests. Social capital has a stronger effect on transfers between research areas that are more cognitively “distant”. Furthermore, institutional stability is associated with more conservative research interests. Overall, we contribute: i) a theoretical account of the features of change in scientists’ research interests; ii) a novel methodological approach that introduces a model of scientists’ trajectories and measures of intellectual capital, social capital, diversity, and power; and finally, iii) empirical evidence from high-energy physics. More generally, this paper addresses the relative lack of empirical works within the body of literature that investigates science as a cultural evolutionary system [12].
@@ 1.2 Conceptual framework
Scientists’ opportunities are constrained by their prior knowledge, since entering new research areas involves learning costs. Yet, the exploitation of prior knowledge may require adjustments to new realities. Through a gradual process of adaptation, scientists can strategically retain the benefits of “problem retention” (e.g., the exploitation of “accumulated skills and resources” in one area, or “of an established research network”, [13, p. 106]) while progressively investing resources in new research directions. This is illustrated in Figure 1, which represents the research agendas of one scientist during two distinct time periods. Cells indicate the resources exploited by the scientist (e.g., concepts, models, methods, or techniques) and colors indicate to what problem areas this knowledge applied. A resource can be applied simultaneously to different ends, as expressed by the hatched cell; however, resources are often tailored for specific research purposes. Figure 1 also shows how scientists can enter new research areas by repurposing their knowledge to new ends [18, 19]. We call this strategy “conversion”, in reference to the typology of incremental institutional change proposed by [20, ]. Indeed, as shown in previous works on the transformations of high-energy physics facilities to photon science instruments [21, 22, 23], historical institutionalism can account for gradual adaptations with large cumulative effects taking place in response to scientific and technological change [24]. In this paper, we apply the typology of change to individuals rather than organizations.
@@
Scientists manage two kinds of assets when navigating the trade-off between specialization and adaptation: their own prior expertise, and the expertise to which they have access through their social network. Both constitute “capital” [17], i.e. assets that individuals (e.g. scientists) can accumulate and leverage in the competitive context of their field. “Capital” (whether “economic”, “cultural”, “social” or even “symbolic”, cf. [17]) defines the scope of scientists’ opportunities and therefore their ability to revise their research agenda. This paper considers the cultural capital possessed by scientists in the form of scientific knowledge (which will be referred to as “intellectual capital”), and social capital. Measures that represent these concepts will be proposed, and their effect on the magnitude of transfers of attention across research areas will be evaluated. Emphasis will be put on the divide underlined in [11] between two dimensions of social capital, namely “power” (in fact, network size in the present paper) and “diversity” (of cognitive resources). Group diversity is generally recognized as a factor of adaptation in an evolving environment [25, 26, 27]. “Power” is also plausibly associated with higher abilities.
@@ 1.3 The case of high-energy physics: navigating a changing epistemic landscape
The field of particle physics has relied on the input of increasingly large particle colliders to achieve progress. These efforts culminated in 2010 with the start of the Large Hadron Collider (LHC), the largest accelerator ever. However, the LHC has found no evidence for anything that was not already predicted by the Standard Model of particle physics, and it is increasingly plausible that no future accelerator could ever find any evidence for new particles, leading to a situation of “crisis” [28]. Although the LHC will continue to take data for years and plans for future accelerators are being discussed [29], some have even speculated that particle physics as we know it has come to an end [30, 31], “the proscenium [being] captured by astrophysics and cosmology” instead [31]. In a recent assessment of the field, physicist Mikhail Shifman claimed that the “pause in accelerator programs we are witnessing now is not necessarily […] the end of explorations at [high energies]”; instead, such explorations “will continue, perhaps in a new form, with novel devices” [32]. Indeed, new experimental opportunities have emerged in parallel. Gravitational Waves astronomy began in 2015, with the first “direct” detection of such waves from a black-hole merger by LIGO [33]. Other opportunities include direct searches of dark matter of astrophysical origin in underground facilities (e.g. LUX and Xenon experiments), and more precise data on the early universe from the Planck observatory. In fact, astrophysics seems to be increasingly replacing particle colliders in citations across experimental and theoretical high-energy physics. [34]. These transformations have unfolded gradually: the LHC started in 2010 but it operated much below its target energy until 2015; Planck released data on the cosmic microwave background between 2010 and 2018; etc. (see Figure 9, Appendix A.1). The present paper investigates how high-energy physicists have adapted in reaction to these changes in their field.
@@ 2.1 Data
Our source is the Inspire HEP database [35]. It aggregates High-Energy Physics (HEP) literature from various sources, including the main scientific publishers and arXiv, and has been used in a number of works [34, 36, 37, 38, 39]. For the literature on HEP, it is more comprehensive than the APS dataset that has been the focus of previous work on the evolution of research interests, because the latter is limited to a few journals [^1] https://journals.aps.org/datasets. The database is free-access and includes semantic data for each paper (mainly their title and abstract, sometimes also keywords and PACS categories), curated authorship and institutional affiliations, lists of references and associated experiments. This database implements both automatized and manual measures for the disambiguation of author names [^2] Besides the use of “advance algorithms” of author-disambiguation, Inspire invites scientists to correct their own publication record on the website (cf. https://twiki.cern.ch/twiki/pub/Inspire/WebHome/INSPIRE_background.pdf, June 2014) , thus allowing careers’ analyses [38]. Nevertheless, occasional misidentifications remain possible.
@@ 2.2 Measuring research agendas
Our strategy for measuring research portfolios relies neither on PACS categories (in contrast to [4, 14, 5, 7]) nor on citation data (unlike [6]). Instead, research areas are extracted with a topic model that recovers latent “topics” within the corpus and their vocabulary distributions [41, 42]. This makes the approach applicable to a wider range of situations for which only textual data (even short texts) are available. Arguably, texts also provide a more direct access to “behavior” or the structure of knowledge. A coarse-grained classification of the literature is performed into $K=20$ broad “topics”. The number of topics is always somewhat arbitrary, but this is not as much an issue of topic models themselves: as [13, ] puts it, “in such analyses [of problem change], empirical findings will in part reflect the defined scope of problem areas”, which is itself arbitrary. In fact, that topic models give some control over the cognitive scale is a feature, rather than a bug [^4] Previous works based on the PACS categories have leveraged the different levels of this hierarchical classification system to investigate different scales..
We use the Embedding Topic Model developed by [43], which relies on pretrained embeddings representations of the keywords thus providing more reliable classifications of heavy-tailed vocabulary distributions. We train the model the abstracts of the $D=186,162$ documents published between 2000 and 2019. Keywords are extracted from the papers’ titles and abstracts by filtering n-grams between one and three words matching syntactic expressions susceptible of carrying scientific information (by designating concepts, models, methods, etc.), following the procedure described in [34, 44]. Keywords’ embeddings are learned using a skip-grap model in $L=50$ dimensions (few are needed given the small size of the vocabulary, $V=4,751$). We obtain the topics listed in Appendix A.3, Table 1. This table shows their most frequent keywords, as well as the PACS categories that are most correlated with them. Four of the 20 resulting topics regroup keywords that do not clearly refer to any specific research area (e.g. “paper”, “approach”) and correlate poorly with the PACS categories, which suggests their lack of scientific dimension. Consequently, the present analysis only considers the $K=16$ remaining topics that designate actual research areas (as confirmed by their strong tendency to cite themselves, cf. Appendix A.3.3, and their correlation with the PACS classification, cf. Appendix A.3.4).
@@ 2.3 Measuring capital
Evaluting the effect of intellectual and social capital on scientists’ trajectories requires quantitative measures of these concepts. An extensive review of prior usage of Bourdieu’s notion of capital in bibliometrics has been provided by [45, p. 198-200 ], with most references addressing symbolic and social capital, and none addressing cultural capital. In the following, a consistent approach to intellectual and social capital is proposed, from which measures of diversity and power are derived.
@@ 2.3.2 Social capital
Many measures of scientists’ social capital have been proposed [11, 45], the simplest being the amount of collaborators of a scientist (i.e. degree centrality in the co-authorship network). Other measures are related to betweeness centrality, which concerns the extent to which an actor “bridges” a network (e.g. “brokerage”, i.e., the ability of an individual to overcome “structural holes” in a social network, [46]). The difference between these measures matter; for instance, redundant contacts (and sheer network size) can be detrimental. [11, ] distinguishes two general approaches to social capital, depending on whether the emphasis is placed on “power” versus “diversity”. Measures of social capital (as those reviewed and proposed in [11]) typically represent social capital by single scalars; however, social capital has multiple dimensions. In fact, according to [16, ], “the volume of social capital possessed by a particular agent […] depends on the extent of the network of links that he can effectively mobilize, and on the volume of capital (economic, cultural or symbolic) possessed by each of those to whom he is linked”. In that respect, social capital can come in different forms depending on the resources being leveraged via one’s network. In the following, we focus on the intellectual dimension of social capital, which we represent by a vector $\bf{S}_{a}$ that encodes the amount of expertise in each research area available to scientists via their network. $\bf{S}_{a}$ is defined as the sum of the intellectual capital of $a$’s collaborators, weighted by the strength of their relationship:
@@ 2.3.3 Diversity and power
Measures of “diversity” (and “power”) can be readily derived from $\bm{I_{a}}$ (and $\bm{S_{a}}$). A common measure of diversity is the Shannon entropy $H$ [^8] The Shannon entropy of the discrete uniform distribution on $\{1,\dots,n\}$ is $\log{n}$, such as that the exponential of the entropy is roughly the effective amount of outcomes for the distribution ($n$).. Let $D(\bm{I_{a}})=\exp{H(\bm{I_{a}})}$ be the diversity of intellectual capital (and $D(\bm{S_{a}})=\exp{H(\bm{S_{a}})}$ that of social capital). Roughly speaking, these are measures of how many research areas scientists have divided their cognitive or social resources among. In the cohort, individuals typically have cognitive resources in several research areas ($\mu_{D(\bm{I_{a}})}=6.1,\sigma=2.2$), and social capital is even more diverse ($\mu_{D(\bm{S_{a}})}=8.5,\sigma=2.2$). In fact, intuitively, scientific collaborations enable individuals to take advantage of their group’s cognitive diversity. Furthermore, $D(\bm{I_{a}})$ and $D(\bm{S_{a}})$ are highly correlated ($R=0.75$); indeed, individuals with more diverse interests are more able to engage with diverse collaborators. Since the diversity of social capital is only expected to enhance individuals’ abilities when it exceeds that of their own intellectual resources, from now on, only excess social capital diversity $D^{\ast}(\bm{S_{a}})$ (defined as the residuals of the linear regression of $D(\bm{S_{a}})$ against $D(\bm{I_{a}})$, by ordinary least squares) is considered999[7, ] noted the difficulty of determining the direction of the causal relationship between social resources and research interests. This approach aims to overcome the very same challenge..
@@
For each of these (diversity and power), an alternative measure is considered for robustness assessment (based on Stirling’s index for diversity and brokerage for power). All measures are compared in Appendix A.4.